# Heeji Yoon1*  Jaewoo Jung1*  Junwan Kim2*Hyungyu Choi3 Heeseong Shin1Sangbeom Lim4Honggyu An1Chaehyun Kim1Jisang Han1 Donghyun Kim4

### v1

@@ 4.3 Visual Representation Alignment for MLLMs
Beyond residual connection, we further explore a more principled approach, which is to explicitly align intermediate visual representations with the encoder features (Yu et al., 2024); see Fig. 2(d). Let $\mathbf{z}$ denote the frozen encoder features from $V_{\psi}(\cdot)$ and $\mathbf{e}^{\mathrm{img}}_{\ell}\!\in\!\mathbb{R}^{N\times D}$ the visual representations at the $\ell$-th layer of the MLLM. We introduce a learnable projection $P_{\pi}$ to map $\mathbf{e}^{\mathrm{img}}_{\ell}$ into the encoder feature space and define the visual representation alignment loss:
@@
As shown in Fig. 2(e,f), this alignment outperforms residual connection both in CKNNA similarity and on multimodal benchmarks. This indicates that directly constraining intermediate features offers a more effective mechanism for preserving fine-grained visual semantics by regularizing the representation space, rather than simply forwarding encoder features, which makes the aligned features easier for the language model to integrate.
Despite the general performance boost from re-injecting projected vision tower features and adopting representation alignment, a notable exception is MMVP (Tong et al., 2024b), which is designed to stress cases where CLIP-like features underperform. In this setting, we observe a small drop in performance, suggesting that naively propagating the vision tower’s features can also transmit its inductive biases and limitations. These findings naturally prompt the question of the alignment target: should the model maintain alignment with the original encoder features $\mathbf{z}$, or can it be guided to preserve more informative visual semantics? While aligning to $\mathbf{z}$ helps retain salient attributes, its utility is bounded by the encoder’s representational capacity.
@@
Figure 3: Illustration of VIRAL. Building upon our findings in visual representation alignment, we align visual pathway representation from MLLMs to strong, informative representations from VFMs to improve the vision understanding performance of MLLMs.
Motivated by this, we adopt stronger vision foundation models (VFMs) as teachers to supervise internal visual representations, providing richer, vision-centric targets that complement language supervision. Building on this insight, we propose VIsual Representation ALignment (VIRAL), which aligns intermediate MLLM visual representations to features from a pretrained VFM, thereby preserving richer visual semantics than those available from the input encoder alone. Let $\mathcal{E}(\cdot)$ denote a pretrained VFM encoder. Given an input image $I$, the encoder produces target features $\mathbf{y}=\mathcal{E}(I)\in\mathbb{R}^{N\times d}$, where $d$ is the VFM channel dimension. Let $\mathbf{e}^{\mathrm{img}}_{\ell}\in\mathbb{R}^{N\times D}$ be the MLLM’s visual representations at layer $\ell$, and let $P_{\pi}$ be a learnable projection that maps $\mathbf{e}^{\mathrm{img}}_{\ell}$ into the VFM feature space. We instantiate the visual representation alignment loss by replacing the encoder target $\mathbf{z}$ in Eq. 5 with $\mathbf{y}$:
@@
Minimizing $\mathcal{L}_{\mathrm{VRA}}$ regularizes the MLLM’s internal visual pathway to align with the VFM, thereby preserving informative spatial and semantic structure. The framework is illustrated in Fig. 3.
@@ 5.1 Experimental Settings
We build on the widely used LLaVA-1.5 (Liu et al., 2024), which couples Vicuna-1.5 (Chiang et al., 2023) as the language model with a CLIP vision encoder (Radford et al., 2021). Following its instruction-tuning recipe, we adopt LoRA (Hu et al., 2022) for efficient adaptation; prior work reports that LLaVA-1.5 with LoRA attains performance comparable to full fine-tuning. Unless otherwise noted, we use only the original LLaVA-665K dataset (Liu et al., 2024) without any additional data. The visual-representation projector $P_{\pi}$ is a lightweight three-layer MLP with SiLU activations. All experiments are run on four NVIDIA A100 (40 GB) GPUs.
To supervise internal visual representations, we employ a diverse set of pretrained vision foundation models (kept frozen, with stop-gradient on their features): DINOv2 (Oquab et al., 2023), CLIP (Radford et al., 2021), Depth Anything V2 (Yang et al., 2024), RADIO-v2.5 (Heinrich et al., 2025), and SAM (Kirillov et al., 2023).
@@
To demonstrate the effectiveness of VIRAL, we evaluate it on widely used benchmarks across three categories: (1) vision-centric tasks requiring spatial reasoning or object counting, including CV-Bench2D (Tong et al., 2024a), What’s Up (Chen et al., 2025b; Kamath et al., 2023), and MMVP (Tong et al., 2024b); (2) multimodal hallucination detection, using POPE (Li et al., 2023); and (3) general multimodal understanding, assessed via MME (Yin et al., 2024), MMStar (Chen et al., 2024b). This selection of benchmarks is motivated by the goals of our method: improving visual grounding should enhance performance on vision-centric and hallucination-sensitive tasks, while ensuring strong performance on general multimodal benchmarks to preserve overall capability. For evaluation, we report overall accuracy for CV-Bench2D, MMVP, What’s Up, POPE, and MMStart and total score for MME, following the respective benchmark protocols.
@@ 5.2 Main Results
60.50% 36.07% 63.57% 84.92% 39.67% 1765.65
Vicuna-1.5-13B CLIP 57.51% 39.33% 44.44% 87.12% 34.47% 1599.04
58.97% 45.33% 62.26% 87.79% 37.00% 1636.62
Table 1: Effect of visual representation alignment. We compare models trained with and without $\mathcal{L}{\text{VRA}}$ across various vision encoders and LLM backbones, evaluating them on both vision-centric and general multimodal benchmarks. Our simple regularization, $\mathcal{L}_{\text{VRA}}$, combined with DINOv2 (Oquab et al., 2023), consistently improves performance across all encoders.
The results on vision-centric benchmarks, visual hallucination tasks, and general vision–language evaluations are summarized in Table 1. Across identical training settings, the model trained with VIRAL consistently outperforms the baseline, with the largest gains on vision-centric tasks that demand fine-grained visual understanding. These improvements arise from a simple intervention—aligning intermediate MLLM features to VFM targets—which effectively enhances the visual pathway.
To test whether these gains merely compensate for the limitations of a contrastive-only vision encoder, we further evaluate with SigLIPv2 (Tschannen et al., 2025), which is trained using both contrastive and visually self-supervised objectives. Even with a much stronger vision encoder, adding our alignment loss yields consistent improvements. Moreover, to examine whether our method follows a scaling trend and is not confined to a particular LLM, we also include results with a scaled-up backbone, comparing Vicuna-1.5-13B against 7B, and with an alternative language backbone, Qwen2.5-7B (Bai et al., 2025). Taken together, these findings highlight a broader principle: regularizing intermediate visual representations is a generally applicable strategy that strengthens MLLMs across vision encoders, model scales, and language backbones.
5.3 Component-wise Analysis
VFM Layer Index Objective CV-Bench2D MMVP What’s Up POPE MME
@@ 5.3 Component-wise Analysis
DINOv2 16 Cos. Sim. 59.67% 33.33% 48.55% 88.32% 1694.52
CLIP 16 Cos. Sim. 57.51% 29.33% 44.50% 87.17% 1548.49
SAM 16 Cos. Sim. 57.58% 30.27% 49.84% 88.34% 1648.77
DAv2 16 Cos. Sim. 58.55% 28.67% 47.29% 88.70% 1682.42
RADIO 16 Cos. Sim. 57.59% 31.80% 47.35% 88.52% 1692.94
Ablation studies on different single-layer targets
DINOv2 4 Cos. Sim. 58.55% 30.67% 45.05% 87.68% 1720.36
DINOv2 8 Cos. Sim. 58.28% 27.70% 48.32% 88.43% 1662.67
DINOv2 12 Cos. Sim. 57.77% 28.59% 48.19% 88.27% 1648.88
DINOv2 16 Cos. Sim. 59.67% 33.33% 48.55% 88.32% 1694.52
DINOv2 20 Cos. Sim. 55.22% 27.41% 48.04% 88.39% 1705.97
DINOv2 24 Cos. Sim. 55.77% 27.48% 47.99% 88.10% 1740.55
DINOv2 28 Cos. Sim. 54.87% 27.19% 47.82% 88.56% 1755.86
DINOv2 32 Cos. Sim. 56.12% 26.52% 47.60% 87.32% 1678.69
Ablation studies on different multi-layer targets
DINOv2 151715-17 Cos. Sim. 59.32% 28.00% 47.17% 87.61% 1639.72
DINOv2 141814-18 Cos. Sim. 49.62% 22.55% 42.58% 87.90% 1444.32
Ablation studies on different alignment objectives
DINOv2 16 Relation 58.83% 26.60% 49.05% 87.58% 1674.30
DINOv2 16 Cos. Sim. 59.67% 33.33% 48.55% 88.32% 1694.52
Table 2: Ablation study on key design components. We analyze the effects of (i) different vision foundation models (VFMs), (ii) alignment target layers, and (iii) alignment objectives, through evaluation on both vision-centric and general multimodal benchmarks. All experiments are conducted on the LLaVA-1.5-7B baseline.
In this ablation study, we conduct a comprehensive analysis of key design choices underlying our framework, focusing on three core components: the selection of target visual features, the choice of alignment layer, and the formulation of alignment objectives. As in Table 2, we evaluate the impact of each component across five benchmarks—CV-Bench, MMVP, What’s Up, POPE, and MME—to elucidate their respective contributions to the model’s performance on vision-grounded tasks.
@@
We begin by identifying the most effective target visual features for enhancing the alignment of internal visual representations within MLLMs. While residual connections and alignment with CLIP (LLaVA’s original encoder) help improve visual comprehension (Figure 2), their performance on spatial tasks like MMVP is limited—likely due to CLIP’s weakness in modeling spatial relations. To address this, we evaluate several stronger vision foundation models (VFMs), including DINOv2, CLIP, Segment Anything, Depth Anythingv2, and RADIO. As shown in Table 2, our analysis confirms that aligning with stronger visual features indeed enhances visual understanding, with DINOv2 and other VFMs demonstrating improved performance compared to CLIP. Results show that DINOv2 consistently emerges as the most effective and versatile, and we thus adopt DINOv2 as the default visual foundation model for all subsequent experiments.
@@
Next, we examine performance of aligning features at individual target layers to identify the most effective alignment position and range. As shown in the single-layer targets ablation results in Table 2, we report performance at every 4th layer throughout the network. We observe that performance varies depending on the alignment layer, with the 16th layer of the 32-layer model consistently yielding stronger results across multiple benchmarks. This trend aligns with prior findings (Zhang et al., 2025b; Kaduri et al., 2025) and our earlier analysis, suggesting that certain intermediate layers in MLLMs are more attuned to visual information processing. To investigate the effective number of target layers, we evaluate multi-layer targets around the 16th—specifically $\pm 1$ (15–17) and $\pm 2$ (14–18) ranges—and observe that applying alignment solely at the 16th layer achieves the best performance. These findings highlight that aligning visual representations at a specific pathway responsible for visual representation processing, rather than uniformly across multiple layers, is more effective in enhancing the visual understanding capabilities of MLLMs. Based on this observation, we adopt the 16th layer as the default alignment target with DINOv2.
Alignment objectives.
We investigate the impact of different feature alignment objectives during instruction tuning. Specifically, we compare the performance of models trained with a feature relation alignment objective, as a substitute for the proposed direct visual representation alignment loss. Here, the alignment objective is defined as a mean squared error (MSE) loss between the self-similarity matrices of the VFM features and the transformed intermediate representations, which effectively distills the structural relationships among visual features following recent approaches (Zhang et al., 2025a; Bolya et al., 2025). As shown in Table 2, we find that cosine similarity-based alignment loss yields higher performance, and adopt it as our default strategy for alignment.
@@ 5.4 Attention Analysis
We analyze the effectiveness of our proposed framework with visual representation alignment in terms of text-to-image attention, as shown in Figure 4. The attention map produced by the $\mathcal{L}_{\mathrm{VRA}}$ trained model exhibits more semantically aligned focus on image regions corresponding to the given textual prompts. We also quantitatively assess attention localization using spatial entropy (Batty, 1974), motivated by the findings of (Kang et al., 2025), which demonstrate that certain attention heads at specific layers are responsible for effectively localizing referred objects, and that such behavior can be identified through spatial entropy. As shown in Figure 4(right), LLaVA-1.5-7B exhibit high entropy across layers and heads, reflecting less discriminative attention patterns. In contrast, our model consistently shows lower entropy—particularly at the aligned intermediate layer where fine-grained visual information is integrated—suggesting that alignment with VFM representations enables more selective and meaningful attention.
5.5 Training Efficiency
$\mathcal{L}_{\mathrm{VRA}}$ 1K 2K 3K 4K 5K
84.2% 87.5% 86.8% 83.7% 84.8%
87.3% 87.9% 87.2% 88.1% 88.0%
(a) POPE
$\mathcal{L}_{\mathrm{VRA}}$ 1K 2K 3K 4K 5K
57.4% 52.8% 54.5% 52.9% 52.8%
48.9% 55.6% 59.9% 58.9% 59.9%
(b) CV-Bench2D
$\mathcal{L}_{\mathrm{VRA}}$ 1K 2K 3K 4K 5K
24.7% 20.0% 20.7% 22.0% 24.0%
20.4% 25.7% 27.1% 30.0% 30.2%
(c) MMVP
Table 3: Training efficiency. Performance with and without $\mathcal{L}_{\mathrm{VRA}}$ evaluated every 1K steps on (a) POPE, (b) CV-Bench2D, and (c) MMVP.
To further assess additional benefits of VIRAL, we evaluate model performance on vision-centric benchmarks, including POPE, CV-Bench2D, and MMVP, at every 1K training steps from the total 5,195 training steps of the visual instruction tuning stage of LLaVA-1.5 in Table 3. Our results show that models trained with VIRAL achieve significantly faster convergence, demonstrating improved performance even in early training stages. This indicates that aligning intermediate visual representations with strong vision foundation models not only enhances final task accuracy but also accelerates training.
@@ 5.6 Robustness Analysis
We further investigate whether our representation alignment loss enables MLLMs to better capture visual information, such as spatial relationships shown in images. Recent observations (Qi et al., 2025) reveal that MLLMs often ignore important visual cues, such as spatial positions, demonstrated by minimal performance degradation even when the order of visual tokens is randomly permuted. To address this, we evaluate whether MLLMs become more sensitive to these random permutations after training with our proposed representation alignment loss.
@@
Table 4: Robustness to token permutation. Number of correct predictions out of 788 spatial reasoning tasks in CV-Bench2D. Models with $\mathcal{L}_{\mathrm{VRA}}$ show larger performance drops under random permutation, indicating stronger sensitivity to spatial relationships.
Specifically, after extracting visual features $\mathbf{z}=V_{\psi}(I)$ from image $I$, we randomly permute the visual tokens before feeding them into the language model $\mathrm{LM}_{\theta}(\cdot)$. We measure performance differences between the original and randomly permuted inputs in the spatial reasoning category of the CV-Bench2D dataset, as shown in Table 4. The results indicate that the original baseline, trained with text-only supervision, exhibits minimal performance drops even when provided with randomly permuted inputs. In contrast, when trained with our proposed visual representation alignment loss, the model experiences a significantly larger performance degradation under random permutation, indicating enhanced sensitivity to spatial token ordering. This observation confirms that our loss successfully encourages MLLMs to better capture and utilize fine-grained spatial relationships in images, rather than disregarding them as irrelevant cues.

### v2

@@ 4.3 Visual Representation Alignment for MLLMs
Beyond residual connection, we further explore a more principled approach, which is to explicitly align intermediate visual representations with the encoder features (Yu et al., 2024); see Figure 2-(c). Let $\mathbf{z}$ denote the frozen encoder features from $V_{\psi}(\cdot)$ and $\mathbf{e}^{\mathrm{img}}_{\ell}\!\in\!\mathbb{R}^{N\times D}$ the visual representations at the $\ell$-th layer of the MLLM. We introduce a learnable projection $P_{\pi}(\cdot)$ to map $\mathbf{e}^{\mathrm{img}}_{\ell}$ into the encoder feature space and define the visual representation alignment loss:
@@
As shown in Figure 2-(d,e), this alignment outperforms residual connection in both CKNNA similarity and multimodal benchmarks. Further analysis on this finding is provided in Appendix B. This shows that constraining intermediate features through an alignment loss offers stronger preservation of fine-grained semantics through explicit regularization, while residual connections offers only weak constraints without enforcing consistency at the feature level.
Despite the general performance boost from retaining encoder-aligned visual information, either by re-injecting projected features or applying visual representation alignment, a notable exception is MMVP (Tong et al., 2024b), which targets cases where CLIP-like features underperform. In this setting, performance shows only marginal improvement or even a slight drop, suggesting that propagating the encoder’s features can also transmit its inductive biases and limitations. These findings raise the question of the alignment target: should the model remain tied to the original encoder features $\mathbf{z}$, or be guided toward more informative visual semantics? While aligning to $\mathbf{z}$ helps retain meaningful attributes, its utility is constrained by the encoder’s representational capacity.
@@
Figure 3: Illustration of VIRAL. We align visual pathway representation from MLLMs to strong, informative representations from VFMs to improve the vision understanding performance of MLLMs.
Motivated by this, we adopt stronger vision foundation models (VFMs) as teachers to supervise internal visual representations, providing richer vision-centric targets that complement language supervision. Building on this insight, we propose VIsual Representation ALignment (VIRAL), which aligns intermediate MLLM visual representations with features from a pretrained VFM, thereby preserving richer visual semantics than those available from the encoder alone. Let $\mathcal{E}(\cdot)$ denote a pretrained VFM encoder. Given an input image $I$, the encoder produces target features $\mathbf{y}=\mathcal{E}(I)\in\mathbb{R}^{N\times d}$, where $d$ is the VFM feature dimension. Let $\mathbf{e}^{\mathrm{img}}_{\ell}\in\mathbb{R}^{N\times D}$ be the MLLM’s visual representations at layer $\ell$, and let $P_{\pi}(\cdot)$ be a learnable projection that maps $\mathbf{e}^{\mathrm{img}}_{\ell}$ into the VFM feature space. We instantiate the visual representation alignment loss by replacing the encoder target $\mathbf{z}$ in Eq. 4 with $\mathbf{y}$:
@@
Minimizing $\mathcal{L}_{\mathrm{VRA}}$ regularizes the MLLM’s internal visual pathway to align with the VFM. The overall framework is illustrated in Figure 3.
@@ 5.1 Experimental Settings
We build on the widely used LLaVA-1.5 (Liu et al., 2024a), leveraging Vicuna-1.5 (Chiang et al., 2023) as the language model with a CLIP vision encoder (Radford et al., 2021). Following its instruction-tuning recipe, we adopt LoRA (Hu et al., 2022) for efficient adaptation as prior work reports that LLaVA-1.5 with LoRA attains comparable performance to full fine-tuning (Liu et al., 2024a). Unless otherwise noted, we use the original LLaVA-665K dataset (Liu et al., 2024a) without any additional data. The visual-representation projector $P_{\pi}(\cdot)$ is a lightweight three-layer MLP with SiLU activations, and we set $\mathcal{E}(\cdot)$ to DINOv2 as default (Section 5.3).
@@
To demonstrate the effectiveness of VIRAL, we evaluate it on widely used benchmarks across three categories: (1) vision-centric tasks requiring spatial reasoning or object counting, including CV-Bench2D (Tong et al., 2024a), What’s Up (Chen et al., 2025b; Kamath et al., 2023), and MMVP (Tong et al., 2024b); (2) multimodal hallucination detection, using POPE (Li et al., 2023b); and (3) general multimodal understanding, assessed via MME (Yin et al., 2024), MMStar (Chen et al., 2024b). These benchmarks align with goals of our method: improving visual grounding should enhance performance on vision-centric and hallucination-sensitive tasks, while ensuring strong performance on general multimodal benchmarks to preserve overall capability. For evaluation, we report overall accuracy for CV-Bench2D, MMVP, What’s Up, POPE, and MMStar and total score for MME. Additional details on the evaluation settings are provided in Appendix A.
@@ 5.2 Main Results
Table 1: Effects of visual representation alignment. We compare models trained with and without $\mathcal{L}_{\text{VRA}}$ across various vision encoders and LLM backbones, evaluating them on both vision-centric and general multimodal benchmarks. Our simple regularization, $\mathcal{L}_{\text{VRA}}$, combined with DINOv2 (Oquab et al., 2023), consistently improves performance across all encoders.
@@
60.50%(+1.53) 36.07%(+2.60) 63.57%(+4.49) 84.92%(-0.96) 39.67%(+0.47) 1765.65(+22.09)
Vicuna-1.5-13B CLIP 57.51% 32.30% 44.44% 87.12% 34.47% 1599.04
58.97%(+1.46) 37.80%(+5.50) 62.26%(+17.82) 87.79%(+0.67) 37.00%(+2.53) 1636.62(+37.58)
The results on vision-centric benchmarks, visual hallucination tasks, and general vision–language evaluations are summarized in Table 1. Across identical training settings, the model trained with VIRAL consistently outperforms the baseline—with the largest gains on fine-grained vision-centric tasks while retaining strong performance on general multimodal benchmarks—through a simple strategy that aligns intermediate MLLM features with VFM targets to strengthen the visual pathway.
To test whether the observed gains arise only when using CLIP as the vision encoder—by compensating for the limitations of a contrastive-only encoder with visually self-supervised features—we further evaluate SigLIPv2 (Tschannen et al., 2025) as the vision encoder, which is trained with both contrastive and self-supervised objectives. Even with this stronger encoder, our alignment loss yields consistent improvements, showing that the gains stem from the alignment itself. Moreover, to examine whether our method follows a scaling trend and is not confined to a particular language model, we also include results with a scaled-up backbone, comparing Vicuna-1.5-13B against 7B, and with an alternative language backbone, Qwen2.5-7B (Bai et al., 2025). Taken together, these findings highlight a broader principle: regularizing intermediate visual representations is a generally applicable strategy that strengthens MLLMs across vision encoders, scales, and language backbones.
Table 2: Ablation study on key design components. We analyze the effects of (i) different vision foundation models (VFMs) and (ii) alignment target layers, through evaluation on vision-centric and general multimodal benchmarks. All experiments are conducted on the LLaVA-1.5-7B baseline.
VFM Layer Index CV-Bench2D MMVP What’s Up POPE MME
@@
DINOv2 16 59.67% 33.33% 48.55% 88.32% 1694.52
CLIP 16 59.53% 29.33% 44.50% 88.10% 1548.49
SAM 16 57.58% 30.27% 49.84% 88.34% 1648.77
DAv2 16 58.55% 28.67% 47.29% 88.70% 1682.42
RADIO 16 57.59% 31.80% 47.35% 88.52% 1692.94
Ablation studies on different target layers
DINOv2 4 58.55% 30.67% 45.05% 87.68% 1720.36
DINOv2 8 58.28% 27.70% 48.32% 88.43% 1662.67
DINOv2 12 57.77% 28.59% 48.19% 88.27% 1648.88
DINOv2 16 59.67% 33.33% 48.55% 88.32% 1694.52
DINOv2 20 55.22% 27.41% 48.04% 88.39% 1705.97
DINOv2 24 55.77% 27.48% 47.99% 88.10% 1740.55
DINOv2 28 54.87% 27.19% 47.82% 88.56% 1755.86
DINOv2 32 56.12% 26.52% 47.60% 87.32% 1678.69
5.3 Component-wise Analysis
In this ablation study, we conduct a comprehensive analysis of key design choices underlying our framework, focusing on core components: the selection of target visual features and the choice of alignment layer. As in Table 2, we evaluate the impact of each component across five benchmarks (CV-Bench, MMVP, What’s Up, POPE, and MME) to validate their respective contributions to the model’s performance on vision-grounded tasks. Additional ablation studies on alignment objectives and target layers are provided in Appendix C.
@@ 5.3 Component-wise Analysis
We begin by identifying the most effective target visual features for enhancing the alignment of internal visual representations within MLLMs. While residual connections and alignment with CLIP (LLaVA-1.5’s original vision encoder) help improve visual comprehension (Figure 2), their performance on spatial tasks like MMVP is limited—likely due to CLIP’s weakness in modeling spatial relations (Yuksekgonul et al., 2022). To address this, we evaluate several stronger vision foundation models (VFMs), including DINOv2 (Oquab et al., 2023), CLIP (Radford et al., 2021), Segment Anything (Kirillov et al., 2023) (SAM), Depth Anything v2 (Yang et al., 2024) (DAv2), and RADIOv2.5 (Heinrich et al., 2025). As shown in Table 2, our analysis confirms that aligning with stronger visual features indeed enhances visual understanding, with DINOv2 and other VFMs demonstrating improved performance compared to CLIP. Results show that DINOv2 consistently emerges as the most effective and versatile, and we thus adopt DINOv2 as the default visual foundation model for all experiments.
@@
We then analyze alignment at individual target layers to determine the most effective position. As shown in the target-layers ablation results in Table 2, we report performance at every 4th layer throughout the network. We observe that performance varies depending on the alignment layer, with the 16th layer of the 32-layer model consistently yielding stronger results across multiple benchmarks. This trend is consistent with prior findings (Zhang et al., 2025b; Kaduri et al., 2025) and our earlier analysis, suggesting that certain intermediate layers in MLLMs are particularly attuned to visual information processing.
@@ 5.4 Attention Analysis
We analyze the effectiveness of our proposed framework with visual representation alignment in terms of text-to-image attention, as shown in Figure 4 (left). The attention map produced by the $\mathcal{L}_{\mathrm{VRA}}$ trained model exhibits more semantically aligned focus on image regions corresponding to the given textual prompts. To quantify this, we adopt spatial entropy (Batty, 1974), motivated by (Kang et al., 2025), as a metric of attention localization. As shown in Figure 4 (right), LLaVA-1.5-7B exhibits high entropy across layers and heads, reflecting dispersed attention patterns, whereas our model shows consistently lower entropy—particularly at the aligned intermediate layer—indicating more selective and meaningful attention patterns.
@@ 5.5 Robustness Analysis
Table 3: Robustness to token permutation. Number of correct predictions out of 788 spatial reasoning tasks in CV-Bench2D.
@@
We investigate whether our representation alignment loss enables MLLMs to better capture spatial relationships. Prior work (Qi et al., 2025) shows that MLLMs often overlook spatial cues, exhibiting only minor performance drops even when the order of visual tokens is randomly permuted (see Appendix A). To assess whether our method makes models more sensitive to such cues, we extract visual features $\mathbf{z}=V_{\psi}(I)$ from an image $I$, randomly permute the tokens, and feed them into the language model $\mathrm{LM}_{\theta}(\cdot)$. We then evaluate performance on the spatial reasoning category of CV-Bench2D. Table 3 shows that while the text-only baseline undergoes little degradation under permutation, our model suffers larger drops, reflecting increased sensitivity to spatial structure. This confirms that our loss encourages MLLMs to capture and exploit fine-grained spatial relationships.